# A Bayes Factor Framework for Unified Parameter Estimation and Hypothesis Testing

### v3

@@ 1 Introduction
Figure 1: Examples of P-value functions and support curves. P-value are two-sided. Bayes factors are oriented in favor of the tested parameter value over a specified alternative hypothesis (i.e., a higher Bayes factor indicates higher support for the parameter value over the alternative).
@@
One approach to address this issue is to report the Bayes factor (Jeffreys, , 1939; Good, , 1958; Kass and Raftery, , 1995), i.e., the updating factor of the prior to posterior odds of two hypotheses. As such, Bayes factors allow data analysts to evaluate the relative evidence for two hypotheses without depending on the prior probabilities of the hypotheses; for example, a Bayes factor can quantify the evidence for the presence or absence of a treatment effect without having to assign prior probabilities to these hypotheses (although one still has to specify a prior for the parameter under the alternative, which is challenging in itself). However, the use of Bayes factors comes at the cost of lacking an overarching concept, such as a P-value function or posterior distribution, that can provide data analyst with a coherent set of point and interval estimates. In practice, data analysts who wish to perform hypothesis testing with Bayes factors but also parameter estimation are therefore faced with a dilemma; they can either supply their Bayes factors with a posterior distribution conditional on one hypothesis being true (e.g., the posterior of a treatment effect, assuming the effect is indeed present), which can lead to contradictory conclusions with the Bayes factor (for examples, see Stone, , 1997; Wagenmakers et al., , 2022), or they can assign prior probabilities to the tested hypotheses and report a posterior averaged over both hypotheses (known as Bayesian model averaging, see e.g., Hoeting et al., , 1999; Campbell and Gustafson, , 2022), but this requires specification of prior probabilities which is highly controversial and the reason why the Bayes factor was reported in the first place rather than the posterior probabilities of the hypotheses.
@@
Our goal is therefore to resolve this dilemma and provide a unified framework for estimation and hypothesis testing based on Bayes factors. The idea is the same as for the P-value function; we consider the Bayes factor as a function of the tested parameter. We then use this support curve to derive point estimates, interval estimates, and Bayes factors (as shown in the right plot in Figure 1), establishing a duality between hypothesis testing and parameter estimation. Our framework builds on the recently proposed Bayesian support interval (Wagenmakers et al., , 2022; Pawel et al., , 2024) and extends it with the novel concept of point estimation based on Bayes factors. We call the resulting estimate the maximum evidence estimate (MEE) – the parameter value that receives the most evidential support from the data over a specified alternative hypothesis. This provides data analysts with a unified framework for statistical inference centred around the Bayes factor.
@@
Approaches related to the support curve have recently been proposed in the physics community under the names of ‘Bayes factor surface’ (Fowlie, , 2024) and ‘K ratio’ (Afzal et al., , 2023). Another method called ‘Bayes factor function’ has recently been proposed by Johnson et al., (2023). In this approach, Bayes factors are viewed as a function of a hyperparameter of the prior under the alternative hypothesis but for a fixed null hypothesis. The Bayes factor function can in some cases be made equivalent to the support curve if the roles of the null and alternative hypothesis are reversed and point mass priors are assigned to the parameter under the alternative, yet this seems rather unnatural from the perspective of both methods. Finally, a similar attempt of reconciling Bayesian interval estimation and hypothesis testing with a focus on the ‘Bayesian evidence value’ has been recently investigated by Kelter, (2022).
@@ 2 Support curve inference
Suppose we observe data $y$ with an assumed distribution with probability density/mass function $p(y\mid\theta,\psi)$ that depends on parameters $\theta\in\Theta$ and $\psi\in\Psi$, with $\theta$ being the focus parameters and $\psi$ being possible nuisance parameters. Consider two hypotheses, the null hypothesis $H_{0}\colon\theta=\theta_{0}$ postulating that $\theta$ takes a certain value $\theta_{0}$ and the alternative hypothesis $H_{1}\colon\theta\neq\theta_{0}$ postulating that $\theta$ takes a different value. A natural measure of relative evidence for the two hypotheses is the Bayes factor (Jeffreys, , 1939; Good, , 1958; Kass and Raftery, , 1995), the data-based updating factor of the prior odds of the hypotheses to their posterior odds
@@
with $p(\theta,\psi\mid H_{1})$ denoting the prior assigned to the parameters under $H_{1}$ and $p(\psi\mid H_{0})$ the prior assigned to the nuisance parameters under $H_{0}$.
@@
As (1a) shows, the Bayes factor represents the data-based core of the Bayesian belief calculus. It remains useful even if one rejects the idea of assigning probabilities to $H_{0}$ and $H_{1}$, since this is not necessary (Goodman, , 1999). The alternative expression of the Bayes factor in equation (1b) shows that this update is dictated by the relative predictive accuracy of the two hypotheses. That is, the posterior odds of the null hypothesis $H_{0}$ increase if it outperforms the competing alternative hypothesis $H_{1}$ in predicting the data $y$, and vice versa (Good, , 1952; Gneiting and Raftery, , 2007). Finally, the last equation (1c) shows how the Bayes factor can be calculated, i.e., by dividing the likelihood of $y$ under the null value $\theta_{0}$ (possibly integrated over the prior of $\psi$ under $H_{0}$) by the likelihood of $y$ integrated over the prior of $\theta$ (and possibly $\psi$) under $H_{1}$. The priors for $\theta$ and $\psi$ may also be point mass priors, in which case the Bayes factor reduces to a likelihood ratio.
@@
The idea now is to consider the Bayes factor (1) as a function of $\theta_{0}$, that is, to vary the tested parameter value (the point null hypothesis $H_{0}\colon\theta=\theta_{0}$) in order to assess the support for different parameter values over the alternative $H_{1}$, see the right plot in Figure 1 for an example. Like the P-value function, this support curve (SC) helps to address cognitive challenges with inferential statistics (Greenland, , 2017). For example, it shifts the focus of inference from testing a single privileged null hypothesis (e.g., the hypothesis that there is no treatment effect) to an entire continuum of hypotheses. By looking at the SC, data analysts can then identify hypotheses that receive equal or even less support from the data than the privileged one; for example, a parameter value indicating a very large treatment effect may receive equal support as the value of no treatment effect (sometimes called ‘counternull’, see Rosenthal and Rubin, , 1994; Bind and Rubin, , 2024).
@@
For one- or two-dimensional focus parameters $\theta$, the SC can be plotted as a curve or surface, respectively, so that the relative support for parameter values can be visually assessed. For higher dimensional focus parameters, this becomes more difficult and the SC may need to be summarized in some way, which we discuss in the following.
@@ 2.1 Support sets
The SC can be used to obtain support sets (Wagenmakers et al., , 2022) which are set-valued estimates for $\theta$ based on inverting the Bayes factor (1) similar to how P-value functions are inverted to obtain confidence sets. Specifically, a support set at support level $k>0$ is defined by
@@
that is, the parameter values for which the Bayes factor indicates at least evidence of level $k$ over the specified alternative. In practice, a $k$ support set (typically an interval) is obtained from ‘cutting’ the SC at $k$ and taking the parameter values above as part of the support set (see the right plot in Figure 1 for illustration). It may happen that for certain choices of $k$ the support set is empty because the data do not constitute relative evidence at that level.
The choice of the support level is arbitrary, just as the choice of the confidence level from a confidence set is. One may, for example, report the support level $k=1$ as it represents the tipping point at which the parameter values begin to be supported over the alternative. Conventions for Bayes factor evidence levels can also be used. For example, based on the convention from Jeffreys, (1961), a support set at level $k=10$ includes the parameter values that receive ‘strong’ relative support from the data, while a $k=1/10$ support set includes the parameter values that are at least not strongly contradicted.
@@ 2.3 Example: Normal mean
Suppose we observe a single observation $y$ assumed to be sampled (at least approximately) from a normal distribution $Y\mid\theta\sim\operatorname{N}(\theta,\sigma^{2})$. Assume that $\sigma^{2}$ is known and we want to conduct inferences regarding $\theta$. This is a simple but frequently encountered scenario, for example, $y$ could be an estimated regression coefficient from a generalized linear model and $\sigma$ its estimated standard error. In the following we will consider an example from RECOVERY collaborative group, (2022). This randomised controlled trial found a reduction in mortality of patients hospitalised with COVID-19 when treated with baricitinib compared to usual care (age-adjusted log hazard ratio $y=-0.14$ with standard error $\sigma=0.064$ estimated with Cox regression). To obtain a Bayes factor for contrasting $H_{0}\colon\theta=\theta_{0}$ against $H_{1}\colon\theta\neq\theta_{0}$ we need to formulate a prior for $\theta$ under the alternative $H_{1}$. We will now discuss three choices with different characteristics shown in Table 1.
Table 1: Support curve $\operatorname{BF}_{01}$, maximum evidence estimate $\hat{\theta}_{\mathrm{ME}}$, evidence value $k_{\mathrm{ME}}$, and $k$ support interval (SI) for a normal mean based on one observation $y$ from $Y\mid\theta\sim\operatorname{N}(\theta,\sigma^{2})$ with known variance $\sigma^{2}$ and for three prior distributions for $\theta$ under the alternative $H_{1}$: A normal prior (left), a normal prior centered around the parameter value of the null hypothesis $\theta_{0}$ (middle), a point prior shifted away from the parameter value of the null hypothesis $\theta_{0}$ by $d>0$ (right).
@@
$\operatorname{BF}_{01}$ $\exp\left[-\frac{1}{2}\left\{\frac{(y-\theta_{0})^{2}}{\sigma^{2}}-\frac{(y-m)^{2}}{\sigma^{2}+v}\right\}\right]\sqrt{1+\frac{v}{\sigma^{2}}}$ $\exp\left[-\frac{1}{2}\left\{\frac{(y-\theta_{0})^{2}}{\sigma^{2}(1+\sigma^{2}/v)}\right\}\right]\sqrt{1+\frac{v}{\sigma^{2}}}$ $\exp\left\{\frac{2d(\theta_{0}-y)+d^{2}}{2\sigma^{2}}\right\}$
$\hat{\theta}_{\mathrm{ME}}$ $y$ $y$ non-existent
$k_{\mathrm{ME}}$ $\exp\left\{\frac{(y-m)^{2}}{2(\sigma^{2}+v)}\right\}\sqrt{1+\frac{v}{\sigma^{2}}}$ $\sqrt{1+\frac{v}{\sigma^{2}}}$ non-existent
$k$ SI $y\pm\sigma\sqrt{\log(1+\frac{v}{\sigma^{2}})+\frac{(y-m)^{2}}{\sigma^{2}+v}-\log k^{2}}$ $y\pm\sigma\sqrt{\left\{\log(1+\frac{v}{\sigma^{2}})-\log k^{2}\right\}(1+\frac{\sigma^{2}}{v})}$ $\left[y+\frac{\sigma^{2}\log k}{d}-\frac{d}{2},\infty\right]$
@@
Perhaps the simplest choice is a prior that does not depend on the parameter value $\theta_{0}$ of the null hypothesis, such as a normal prior with mean $m$ and variance $v$ (left column in Table 1). The hyperparameters $m$ and $v$ may be specified based on external data or based on an alternative hypothesis of interest (e.g., the prior mean $m$ could be set to a minimum clinically important treatment effect and $v$ could be set to zero to obtain a point prior as typically used in a power analysis). RECOVERY collaborative group, (2022) reported a meta-analytic log hazard ratio and standard error based on eight previous trials, which we now use to set the prior mean and variance to $m=-0.56$ and $v=0.12^{2}$, see Figure 2 for the resulting SC (orange). In this case, the MEE is given by $\hat{\theta}_{\mathrm{ME}}=y=-0.14$ with the support interval centered around it. Due to the apparent conflict between the observed data and the specified prior under the alternative, the $k=1$ support interval spans an wide range from $-0.35$ to $0.08$, indicating that very beneficial up to slightly harmful treatment effects are supported by the data over the alternative.
@@
The formulae in Table 1 (left column) show that as the prior mean $m$ becomes closer to the observed data $y$, the evidence level $k_{\mathrm{ME}}$ decreases and the support interval becomes narrower. This is because an alternative closer to the data clearly has better predictive accuracy of the data than an alternative further away, and thus fewer point null hypotheses can outpredict it. Figure 2 illustrates this phenomenon with another prior distributions (one with mean at the observed log hazard ratio $y=-0.14$, the blue SC). The SC has its mode still at the observed log hazard ratio estimate but shows a far narrower $k=1$ support interval from $-0.22$ to $-0.06$ than the orange SC with the mean $m=-0.56$ based on the eight previous trials.
@@
Figure 2: Support curve, MEE, and $k=1$ support interval for a log hazard ratio $\theta$ based on estimated log hazard ratio $y=-0.14$ with standard error $\sigma=0.064$ from the RECOVERY trial (RECOVERY collaborative group, , 2022) for different prior distributions for the $\theta$ under the alternative $H_{1}$. A normal likelihood $Y\mid\theta\sim\operatorname{N}(\theta,\sigma^{2}_{i})$ is assumed for the data.
@@
Another approach to formulating a prior distribution for $\theta$ under the alternative commonly suggested in ‘objective’ Bayes theories is to center the prior around the tested parameter value $\theta_{0}$ (Jeffreys, , 1961; Berger and Delampady, , 1987; Kass and Wasserman, , 1995). For example, one can specify a normal prior with mean at the null value $\theta_{0}$ (middle column in Table 1). Thus, the resulting SC varies both the null and the alternative, unlike the SC based on the ‘global’ normal prior with fixed mean $m$. As a result, the interpretation of the SC is different: For such a ‘local’ normal prior, the SC quantifies the support of parameter values over alternative parameter values in a neighborhood around them. As for the global normal prior, the MEE based on the local normal prior is given by $\hat{\theta}_{\mathrm{ME}}=y$ and support intervals are centered around it, but the associated, Bayes factor, evidence level and support interval are different. Figure 2 illustrates that when the mean $m$ of a global normal prior is too different from the observed data $y$ (as in the case of the orange SC, where the prior was specified based on the eight previous trials), the $k=1$ support interval based on the local prior with the same variance is narrower. On the other hand, when the mean $m$ of the global prior is equal to the data (blue SC), the support interval based on the local prior is wider.
@@
The last prior in the right most column of Table 1 represents a point prior shifted from the null value $\theta_{0}$ by $d>0$. The prior is again ‘local’ in the sense that it is different for each tested parameter value of the null hypothesis $\theta_{0}$, and as such encodes an alternative hypothesis that the log hazard ratio is greater than the tested parameter value. However, this leads to an ever-increasing SC, see Figure 2 for a numerical illustration. As a result, the MEE and its evidence level do not exist, while the support interval still exists but its right limit extends to infinity. Although such a prior seems unrealistic, the example demonstrates that a poorly chosen prior can lead to pathological behavior of the resulting SC.

### v4

@@ 1 Introduction
Figure 1: Examples of P-value functions and support curves. P-value are two-sided. Bayes factors are oriented in favour of the tested parameter value over a specified alternative hypothesis (i.e., a higher Bayes factor indicates higher support for the parameter value over the alternative).
@@
One approach to address this issue is to report the Bayes factor (Jeffreys, , 1939; Good, , 1958; Kass and Raftery, , 1995), i.e., the updating factor of the prior to posterior odds of two hypotheses. As such, Bayes factors allow data analysts to evaluate the relative evidence for two hypotheses without depending on the prior probabilities of the hypotheses; for example, a Bayes factor can quantify the evidence for the presence or absence of a treatment effect without having to assign prior probabilities to these hypotheses (although one still has to specify a prior for the parameter under the alternative, which is challenging in itself). However, the use of Bayes factors comes at the cost of lacking an overarching concept, such as a P-value function or posterior distribution, that can provide data analyst with a coherent set of point and interval estimates. In practice, data analysts who wish to perform hypothesis testing with Bayes factors but also parameter estimation are therefore faced with a dilemma; they can either supply their Bayes factors with a posterior distribution conditional on one hypothesis being true (e.g., the posterior of a treatment effect, assuming the effect is indeed present), which can lead to contradictory conclusions with the Bayes factor (for examples, see Stone, , 1997; Wagenmakers et al., , 2022; Kelter, , 2022), or they can assign prior probabilities to the tested hypotheses and report a posterior averaged over both hypotheses (known as Bayesian model averaging, see e.g., Hoeting et al., , 1999; Rouder et al., , 2018; Campbell and Gustafson, , 2022), but this requires specification of prior probabilities which is highly controversial and usually the reason why the Bayes factor was reported in the first place rather than the posterior probabilities of the hypotheses.
@@
Our goal is therefore to resolve this dilemma and provide a unified framework for estimation and hypothesis testing based on Bayes factors. The idea is the same as for the P-value function; we consider the Bayes factor as a function of the tested parameter. We then use this support curve to derive point estimates, interval estimates, and Bayes factors (as shown in the right plot in Figure 1), establishing a duality between hypothesis testing and parameter estimation. This provides data analysts with a unified framework for statistical inference centred around the Bayes factor.
@@
Interval estimates based on the Bayes factor have been proposed under the name of the Bayesian support interval (Wagenmakers et al., , 2022; Pawel et al., , 2024). Similarly, looking at the Bayes factor as a function of parameter values has been proposed in the physics community under the names of ‘Bayes factor surface’ (Fowlie, , 2024) and ‘K ratio’ (Afzal et al., , 2023), and by Rouder and Morey, (2018) in the context of teaching Bayesian statistics. A method called ‘Bayes factor function’ has been proposed by Johnson et al., (2023). In this approach, Bayes factors are viewed as a function of a hyperparameter of the prior under the alternative hypothesis but for a fixed null hypothesis. The Bayes factor function can in some cases be made equivalent to the support curve if the roles of the null and alternative hypothesis are reversed and point mass priors are assigned to the parameter under the alternative, yet this seems rather unnatural from the perspective of both methods. Finally, a similar attempt of reconciling Bayesian interval estimation and hypothesis testing with a focus on the ‘Bayesian evidence value’ has been investigated by Kelter, (2022).
While many elements of support curve inference have been considered in earlier articles, this paper provides a synthesis of these ideas from the perspective of the support curve as a counterpart to the P-value function. A novel contribution is the concept of point estimation based on Bayes factors. We call the resulting estimate the maximum evidence estimate (MEE). The MEE is the parameter value that maximizes the support curve, and as such, is the parameter value that receives the most evidential support from the data over a specified alternative hypothesis. Finally, this paper shows how support curves can be applied in several areas where they have never been used before (meta-analysis, replication studies, and logistic regression).
It is important to note that our intent is not to fuel the ‘statistics wars’ and entertain the position that Bayesian inference or Bayes factors are the best or only valid form of statistical inference. The goal is simply to explore the idea of a Bayes factor analogue of the P-value function, find out how such an approach can be useful, and how it connects to other approaches. Bayes factors, just as P-values and posterior probabilitites, can be delicate tools and users should be wary of their limitations, misuses, and misinterpretations (Tendeiro and Kiers, , 2019; Wong et al., , 2022; Tendeiro et al., , 2024).
@@ 2 Support curve inference
Suppose we observe data $y$ with an assumed distribution with probability density/mass function $p(y\mid\theta,\psi)$ that depends on parameters $\theta\in\Theta$ and $\psi\in\Psi$, with $\theta$ being the focus parameters (the parameters of substantial interest) and $\psi$ being possible nuisance parameters. Consider two hypotheses, the null hypothesis $H_{0}\colon\theta=\theta_{0}$ postulating that $\theta$ takes a certain value $\theta_{0}$ and the alternative hypothesis $H_{1}\colon\theta\neq\theta_{0}$ postulating that $\theta$ takes a different value. A natural measure of relative evidence for the two hypotheses is the Bayes factor (Jeffreys, , 1939; Good, , 1958; Kass and Raftery, , 1995), the data-based updating factor of the prior odds of the hypotheses to their posterior odds
@@
with $p(\theta,\psi\mid H_{1})$ denoting the prior assigned to the parameters under $H_{1}$ and $p(\psi\mid H_{0})$ the prior assigned to the nuisance parameters under $H_{0}$. Note that Bayes factors do not require the null and alternative models to be nested but are applicable to general model comparisons. However, due to their more intuitive interpretation and computational advantages, we will focus on nested models and discuss extensions to non-nested models in Section 5.
@@
As (1a) shows, the Bayes factor represents the data-based core of the Bayesian belief calculus since it transforms the prior odds of $H_{0}$ versus $H_{1}$ to the corresponding posterior odds (Goodman, , 1999). The alternative expression of the Bayes factor in equation (1b) shows that this update is dictated by the relative predictive accuracy of the two hypotheses. That is, the posterior odds of the null hypothesis $H_{0}$ increase if it outperforms the competing alternative hypothesis $H_{1}$ in predicting the data $y$, and vice versa (Good, , 1952; Gneiting and Raftery, , 2007). Since the Bayes factor remains the same for any pair of prior hypothesis probabilities $p(H_{0})$ and $p(H_{1})$, it remains useful to different analysts with different prior hypothesis probabilities, or even if one rejects the idea of assigning probabilities to $H_{0}$ and $H_{1}$. However, this does not mean that the Bayes factor does not depend on any prior distributions; in fact, it can be quite sensitive to the priors assigned to the parameters $\psi$ (and $\theta$) under $H_{0}$ and $H_{1}$, respectively (Kass and Raftery, , 1995). We will return to the choice of parameter priors in Section 2.4. Finally, the last equation (1c) shows how the Bayes factor can be calculated, i.e., by dividing the likelihood of $y$ under the null value $\theta_{0}$ (possibly integrated over the prior of $\psi$ under $H_{0}$) by the likelihood of $y$ integrated over the prior of $\theta$ (and possibly $\psi$) under $H_{1}$. The priors for $\theta$ and $\psi$ may also be point mass priors, in which case the Bayes factor reduces to a likelihood ratio.
@@
The idea now is to consider the Bayes factor (1) as a function of $\theta_{0}$, that is, to vary the tested parameter value (the null hypothesis $H_{0}\colon\theta=\theta_{0}$) in order to assess the support for different parameter values over the alternative $H_{1}$, see the right plot in Figure 1 for an example. Like the P-value function, this support curve (SC) helps to address cognitive challenges with inferential statistics (Greenland, , 2017). For example, it shifts the focus of inference from testing a single privileged null hypothesis (e.g., the hypothesis that there is no treatment effect) to an entire continuum of hypotheses. By looking at the SC, data analysts can then identify hypotheses that receive equal or even less support from the data than the privileged one; for example, a parameter value indicating a very large treatment effect may receive equal support as the value of no treatment effect (sometimes called ‘counternull’, see Rosenthal and Rubin, , 1994; Bind and Rubin, , 2024).
@@
For one- or two-dimensional focus parameters $\theta$, the SC can be plotted as a curve or surface, respectively, so that the relative support for parameter values can be visually assessed. For higher dimensional focus parameters (for example, more than two regression coefficients), this becomes more difficult. In this case, some focus parameters can be re-considered as nuisance parameters to reduce the SC visualization again to a curve or surface. This is similar to ordinary Bayesian inference where a marginal posterior is obtained from marginalizing a joint posterior.
@@ 2.1 Support sets
Apart from visual inspections, we may also want to summarize the SC numerically. The SC can be used to obtain support sets (Wagenmakers et al., , 2022) which are set-valued estimates for $\theta$ based on inverting the Bayes factor (1) similar to how P-value functions are inverted to obtain confidence sets. Specifically, a support set at support level $k>0$ is defined by
@@
that is, the parameter values for which the Bayes factor indicates evidence of at least level $k$ over the specified alternative. A $k$ support set thus contains all the parameter values that, when considered as individual null hypotheses compared to an alternative, receive a Bayes factor of at least $k$. Each included parameter value is therefore associated with an updating factor of at least $k$ or, equally, relative predictive performance of at least $k$. Support sets are thus ‘calibrated’ with respect to prior-posterior updating, as opposed to the calibration of confidence sets with respect to long-run coverage, or that of posterior credible intervals with respect to posterior belief.
In practice, a $k$ support set (typically an interval) is obtained from ‘cutting’ the SC at $k$ and taking the parameter values above as part of the support set (see the right plot in Figure 1 for illustration). It may happen that for certain choices of $k$ the support set is empty because the data do not constitute relative evidence at that level. The choice of the support level is arbitrary, just as the choice of the confidence level from a confidence set is. One may, for example, report the support level $k=1$ as it represents the tipping point at which the parameter values begin to be supported over the alternative. Conventions for Bayes factor evidence levels can also be used. For example, based on the convention from Jeffreys, (1961), a support set at level $k=10$ includes the parameter values that receive ‘strong’ relative support from the data, while a $k=1/10$ support set includes the parameter values that are at least not strongly contradicted.
@@ 2.3 Example: Normal mean
Suppose we observe a mean $\bar{y}$ from $n$ observations, assumed to be sampled (at least approximately) from a normal distribution $\bar{Y}\mid\theta\sim\operatorname{N}(\theta,\kappa^{2}/n)$. Assume that the variance $\kappa^{2}$ is known and we want to conduct inferences regarding $\theta$. For example, suppose we observed a mean test score of $\bar{y}=110$ in a sample of $n=30$ students. Also assume that in the general population the average score is 100 and the standard deviation is $\kappa=15$. To obtain a Bayes factor for contrasting $H_{0}\colon\theta=\theta_{0}$ against $H_{1}\colon\theta\neq\theta_{0}$ we need to formulate a prior for $\theta$ under the alternative $H_{1}$. We will now discuss three choices with different characteristics shown in Table 1.
Table 1: Support curve $\operatorname{BF}_{01}$, maximum evidence estimate $\hat{\theta}_{\mathrm{ME}}$, evidence value $k_{\mathrm{ME}}$, and $k$ support interval (SI) for a normal mean based on an observed mean $\bar{y}$ from $\bar{Y}\mid\theta\sim\operatorname{N}(\theta,\kappa^{2}/n)$ with known variance $\kappa^{2}$ and for three prior distributions for $\theta$ under the alternative $H_{1}$: A normal prior (left), a normal prior centered around the parameter value of the null hypothesis $\theta_{0}$ (middle), a point prior shifted away from the parameter value of the null hypothesis $\theta_{0}$ by $d>0$ (right).
@@
$\operatorname{BF}_{01}$ $\exp\left[-\frac{1}{2}\left\{\frac{(\bar{y}-\theta_{0})^{2}}{\kappa^{2}/n}-\frac{(\bar{y}-m)^{2}}{v+\kappa^{2}/n}\right\}\right]\sqrt{1+\frac{v\,n}{\kappa^{2}}}$ $\exp\left[-\frac{1}{2}\left\{\frac{n(\bar{y}-\theta_{0})^{2}}{\kappa^{2}\{1+\kappa^{2}/(v\,n)\}}\right\}\right]\sqrt{1+\frac{v\,n}{\kappa^{2}}}$ $\exp\left\{\frac{2d(\theta_{0}-\bar{y})+d^{2}}{2\kappa^{2}/n}\right\}$
$\hat{\theta}_{\mathrm{ME}}$ $\bar{y}$ $\bar{y}$ non-existent
$k_{\mathrm{ME}}$ $\exp\left\{\frac{(\bar{y}-m)^{2}}{2(v+\kappa^{2}/n)}\right\}\sqrt{1+\frac{v\,n}{\kappa^{2}}}$ $\sqrt{1+\frac{v\,n}{\kappa^{2}}}$ non-existent
$k$ SI $\bar{y}\pm\frac{\kappa}{\sqrt{n}}\sqrt{\log(1+\frac{v\,n}{\kappa^{2}})+\frac{(\bar{y}-m)^{2}}{v+\kappa^{2}/n}-\log k^{2}}$ $\bar{y}\pm\frac{\kappa}{\sqrt{n}}\sqrt{\left\{\log(1+\frac{v\,n}{\kappa^{2}})-\log k^{2}\right\}(1+\frac{\kappa^{2}}{v\,n})}$ $\bigg{[}\bar{y}+\frac{\kappa^{2}\log k}{n\,d}-\frac{d}{2},\infty\bigg{)}$
@@
Perhaps the simplest choice is a prior that does not depend on the parameter value $\theta_{0}$ of the null hypothesis, such as a normal prior with mean $m$ and variance $v$ (left column in Table 1). The hyperparameters $m$ and $v$ may be specified based on external data or based on an alternative hypothesis of interest. For example, suppose that $50$ students from the previous year had a mean test score of $120$, which we now use to set the prior mean and variance to $m=120$ and $v=15^{2}/50$, see Figure 2 for the resulting SC (orange). In this case, the MEE is given by $\hat{\theta}_{\mathrm{ME}}=\bar{y}=110$ with the support interval centered around it. Due to the apparent conflict between the observed data and the specified prior under the alternative, the $k=1$ support interval spans a wide range from $101.9$ to $118.1$, indicating that average up to above average test scores are supported by the data over the alternative.
@@
The formulae in Table 1 (left column) show that as the prior mean $m$ becomes closer to the observed data $y$, the evidence level $k_{\mathrm{ME}}$ decreases and the support interval becomes narrower. This is because an alternative closer to the data clearly has better predictive accuracy of the data than an alternative further away, and thus fewer point null hypotheses can outpredict it. Figure 2 illustrates this phenomenon with another prior distributions (one with mean at the observed test score $\bar{y}=110$, the blue SC). The SC has its mode still at the observed mean test score but shows a far narrower $k=1$ support interval from $108.1$ to $111.9$ than the orange SC with the mean $m=120$ based on the previous year students.
@@
Figure 2: Support curve, MEE, and $k=1$ support interval for a mean test score $\theta$ based on the observed mean $\bar{y}=110$ from $n=30$ observations for different prior distributions for $\theta$ under the alternative $H_{1}$. A normal likelihood $\bar{Y}\mid\theta\sim\operatorname{N}(\theta,15^{2}/n)$ is assumed for the data.
@@
Another approach to formulating a prior distribution for $\theta$ under the alternative commonly suggested in ‘objective’ Bayes theories is to center the prior around the tested parameter value $\theta_{0}$ (Jeffreys, , 1961; Berger and Delampady, , 1987; Kass and Wasserman, , 1995). For example, one can specify a normal prior with mean at the null value $\theta_{0}$ (middle column in Table 1). Thus, the resulting SC varies both the null and the alternative, unlike the SC based on the ‘global’ normal prior with fixed mean $m$. As a result, the interpretation of the SC is different: For such a ‘local’ normal prior, the SC quantifies the support of parameter values over alternative parameter values in a neighborhood around them. As for the global normal prior, the MEE based on the local normal prior is given by $\hat{\theta}_{\mathrm{ME}}=y$ and support intervals are centered around it, but the associated, Bayes factor, evidence level and support interval are different. Figure 2 illustrates that when the mean $m$ of a global normal prior is too different from the observed mean $\bar{y}$ (as in the case of the orange SC, where the prior was specified based on the students from the previous year), the $k=1$ support interval based on the local prior with the same variance is narrower. On the other hand, when the mean $m$ of the global prior is equal to the data (blue SC), the support interval based on the local prior is wider.
@@
The last prior in the right most column of Table 1 represents a point prior shifted from the null value $\theta_{0}$ by $d>0$. The prior is again ‘local’ in the sense that it is different for each tested parameter value of the null hypothesis $\theta_{0}$, and as such encodes the alternative hypothesis that the mean test score is greater than the tested parameter value. However, this leads to an ever-increasing SC, see Figure 2 for an illustration. As a result, the MEE and its evidence level do not exist, while the support interval still exists but its right limit extends to infinity. Although such a prior seems unrealistic, the example demonstrates that a poorly chosen prior can lead to pathological behavior of the resulting SC.